# Teaching Humans Subtle Differences with DIFFusion

### v1

@@ Abstract
Human expertise depends on the ability to recognize subtle visual differences, such as distinguishing diseases, species, or celestial phenomena. We propose a new method to teach novices how to differentiate between nuanced categories in specialized domains. Our method uses generative models to visualize the minimal change in features to transition between classes, i.e., counterfactuals, and performs well even in domains where data is sparse, examples are unpaired, and category boundaries are not easily explained by text. By manipulating the conditioning space of diffusion models, our proposed method DIFFusion disentangles category structure from instance identity, enabling high-fidelity synthesis even in challenging domains. Experiments across six domains show accurate transitions even with limited and unpaired examples across categories. User studies confirm that our generated counterfactuals outperform unpaired examples in teaching perceptual expertise, showing the potential of generative models for specialized visual learning.
@@ 1 Introduction
Generative models, especially large-scale image diffusion models, have transformed text-to-image creation, opening new ways to visualize concepts across various domains. These models can produce counterfactual images—edits that shift an image from one category to another while preserving its identity—through text-guided editing or classifier-based techniques. To date, such efforts have centered on everyday domains with clear category distinctions, like cats versus dogs, rather than scientific fields such as medical imaging or astronomical classification. In these specialized areas, subtle visual differences are harder to detect, making counterfactuals particularly valuable for enhancing visual learning.
However, applying counterfactuals to specialized domains presents significant challenges. While effective in familiar contexts, counterfactuals struggle in scientific fields where training data is scarce and lacks paired examples showing corresponding features across categories. Text-based editing often fails to capture the precise visual nuances critical in these domains. These challenges are compounded by practical limitations—collecting comprehensive datasets requires specialized expertise and access to restricted equipment or subjects. This combination of limited data representation, unpaired examples, and complex domain-specific variations necessitates novel approaches for effectively leveraging generative models as teaching tools in specialized fields.
@@
Current approaches highlight these limitations. Text-guided editing methods rely on linguistic descriptions, which can be too ambiguous to specify desired visual changes. Methods like Concept Sliders’ [15] effectiveness depend on paired examples in most cases—a constraint limiting their use in teaching scenarios. Visual counterfactual generation methods often require training a classifier, a limitation when data is scarce. Even classifier-free alternatives, like TIME [25], struggle with image quality and coherence for subtle differences.
We propose a framework that teaches visual concepts by generating natural transitions between categories, using class image distributions. Our approach leverages diffusion models’ priors to edit images, smoothly interpolating between categories while preserving instance-specific features without paired data or linguistic descriptions. We achieve this via targeted manipulations in the conditioning space, capturing nuanced transitions and image structure. This allows meaningful transformations from few examples per category, fitting specialized domains with limited data.
Through experiments across six domains, we demonstrate our approach’s effectiveness and utility. We highlight results on medical imaging (retinal diseases) [28], astronomical data (black holes), and natural domains (butterflies), each posing unique challenges in visual expertise training. Our method achieves better reconstruction quality and more coherent transformations compared to existing approaches, with high-quality results from limited data. User studies confirm that our transformations help people distinguish categories more effectively. By generating nuanced counterfactuals in data-scarce specialized domains, our framework bridges the gap between generative modeling and visual expertise training, paving the way for broader applications in scientific education and discovery.

### v2

@@ Abstract
Scientific expertise often requires recognizing subtle visual differences that remain challenging to articulate even for domain experts. We present a system that leverages generative models to automatically discover and visualize minimal discriminative features between categories while preserving instance identity. Our method generates counterfactual visualizations with subtle, targeted transformations between classes, performing well even in domains where data is sparse, examples are unpaired, and category boundaries resist verbal description. Experiments across six domains, including black hole simulations, butterfly taxonomy, and medical imaging, demonstrate accurate transitions with limited training data, highlighting both established discriminative features and novel subtle distinctions that measurably improved category differentiation. User studies confirm our generated counterfactuals significantly outperform traditional approaches in teaching humans to correctly differentiate between fine-grained classes, showing the potential of generative models to advance visual learning and scientific research.
@@ 1 Introduction
Generative models, especially large-scale image diffusion models, have transformed text-to-image creation, opening new ways to visualize concepts across various domains. While these models excel in everyday contexts with clear category distinctions, a far more challenging frontier exists in scientific fields where visual differences between categories are so subtle that they often remain unknown and unidentified even to domain experts.
In specialized scientific domains, the complete set of visual features distinguishing between categories may be partially or entirely undiscovered. For example, astronomers studying black hole simulations have no established verbal characteristics to differentiate MAD from SANE models because these distinguishing features have not yet been comprehensively identified. Entomologists may differentiate Viceroy and Monarch butterflies through the Viceroy’s characteristic cross-sectional black line, yet may miss other distinguishing features that could further help the differentiation. This represents the fundamental challenge for visual expertise training: how do we teach recognition of patterns we ourselves don’t fully understand?
One of the most effective ways to reveal subtle category differences is to transform an image and rapidly flip between the original and its altered version. In scientific domains, this approach faces three key challenges: (1) automatically identifying discriminative features that may not be known or easily articulated even by experts, (2) limiting changes exclusively to these category-defining features, and (3) preserving all other identity characteristics of the instance. We develop a system that combines state-of-the-art image editing techniques with visual algebraic conditioning guidance to address these challenges in data-scarce scientific domains. Our approach automatically identifies discriminative features through visual algebraic operations that extract category-specific information without requiring explicit articulation. By integrating inverted noise maps (z) to preserve identity features with conditioning vectors (c) that guide category transformations, our system achieves effective identity-preserving yet category-changing results, that isolate and visualize subtle differences between scientific categories.
@@
Our approach overcomes limitations in current counterfactual visualization methods, which have traditionally been applied in domains where category distinctions are already well-understood and easily verbalized. Text-guided editing methods rely on linguistic descriptions, which can be too ambiguous to specify desired visual changes. Methods like Concept Sliders’  [15] effectiveness, which is guided by the image distributions themselves, depend on paired examples in most cases—a constraint limiting their use in teaching scenarios. Visual counterfactual generation methods often rely on gradients from a classifier, a limitation when data is scarce. Classifier-free alternatives, like TIME [25], struggle with image quality and coherence for subtle differences.
Through experiments across six domains, we demonstrate our approach’s effectiveness in highlighting visual differences between categories. For instance, in black hole simulations, where distinguishing characteristics between MAD and SANE models remain largely unknown, our counterfactual visualizations emphasize distinct visual patterns in the image distribution. The transformations draw attention to variations in the uniformity of wisps and prominence of the photon ring, which are features that black hole experts themselves had not identified.
User studies confirm the effectiveness of our approach: participants who trained with our counterfactual visualizations demonstrated significantly better category differentiation performance than those using traditional approaches with unpaired images. This validates that our method highlights meaningful visual patterns that can be used to build expertise, even when those subtle patterns have not yet been explicitly identified or understood.